OpenStudy (anonymous):

Please show me how to do this! find all SOLUTIONS of the following equation in the interval [0,2π) cosx=cos3x

OpenStudy (anonymous):

\[\cos x=\cos 3x\] cosx-cos3x=0 \[\cos C-\cos D=2\sin \frac{ C+D }{ 2 }\sin \frac{ D-C }{ 2 }\] use and solve

OpenStudy (anonymous):

\[2\sin \frac{ x+3x }{ 2 }\sin \frac{ 3x-x }{ 2 }=0\] \[\sin 2x \sin x=0\] \[either \sin 2x=0=\sin n \pi \] \[2x=n \pi,n=0,1,2,3,4\] \[x=\frac{ n \pi }{ 2 }\] \[x=0,\frac{ \pi }{ 2 },\pi,\frac{ 3\pi }{ 2 },2\pi \] \[or \sin x=0=\sin n \pi ,x=n \pi,n=0,1,2\] \[x=0,\pi,2\pi\] \[hen ce x=0,\frac{ \pi }{ 2 },\pi,\frac{ 3\pi }{ 2 },2\pi \]